OpenStudy (anonymous):

Find all of the solutions in interval (0, 2π) 2sin ²x = sinx (Also, I'm still not very clear on what an interval is, mind explaining that as well as how to solve this please?)

OpenStudy (anonymous):

it means between x=0 and x=\(\large 2 \pi \), how many times does 2\(sin^2x \) and sinx cross each other. solve this like you would solve this algebra problem: \(\large 2x^2=x \), how would you solve this?

OpenStudy (freckles):

An interval is a set of numbers. For example the interval (0,1] means absolutely all real numbers between 0 and 1 including 1. Example, .055555423 is in that set. So is and irrational number like sqrt(. 5)or even. 5 is in that set. There are in an infinite amount of numbers between 0 and 1.

OpenStudy (anonymous):

\[2x ^{2} = x\] \[x ^{2} = \frac{ x }{ 2 }\]

OpenStudy (anonymous):

ahhhh...I'm not too sure

OpenStudy (freckles):

2x^2-x=0 You should be able to take it from here

OpenStudy (anonymous):

something that adds up to 0 but multiplies to -2 with a factors/sum of factors chart, right? i'm at a stump there, give me another minute to figure it out please

OpenStudy (anonymous):

I'm at a brain fart

OpenStudy (freckles):

Do you know how to factor?